OpenStudy (anonymous):

2a +3ba^2 + 2ab^2 +b^3 how can I get (a+b) (2a^2 +ab + b^2)?

OpenStudy (amistre64):

you seem to be missing something

OpenStudy (anonymous):

what am I missing?

OpenStudy (amistre64):

well, your factored form has 2a^3, but you have no 2a^3 to start with

OpenStudy (anonymous):

it has 2a^2

OpenStudy (amistre64):

a*2a^2 = 2a^3 which isnt in your starting stuff

OpenStudy (anonymous):

thats because its factorized. If I combine the factors I get above but I am not sure how to get to the factor part

OpenStudy (amistre64):

im saying, when you combine your factors, you get 2a^3 ... you have no 2a^3 to start with so one doesnt form into the other

OpenStudy (amistre64):

you are missing something

OpenStudy (anonymous):

ooo sorry my bad, it suppose to be 2a^3 the 2a

OpenStudy (amistre64):

if factor by grouping is your cup o tea, then we want to split it into 6 terms, but grouping is not what im good at

OpenStudy (anonymous):

I'm trying to do that but I am not getting the same as the factor

OpenStudy (anonymous):

im att a(2a^2+2b^2) + b(3a^2 + b^2) which is near close to the factor

OpenStudy (amistre64):

2a^3 +3ba^2 + 2ab^2 +b^3 2a^3 +(2ba^2 + ba^2) + (i ab^2+j ab^2) +b^3 (2a^3 +2ba^2 + i ab^2) + (ba^2+j ab^2 +b^3) a(2a^2 +2ba + i b^2) + b(a^2+j ab +b^2) i loathe by grouping ....

OpenStudy (amistre64):

2a^3 +3ba^2 + 2ab^2 +b^3 2a^3 +(x ba^2 + y ba^2) + (i ab^2 +j ab^2) +b^3 (2a^3 +x ba^2+ i ab^2) + (y ba^2 +j ab^2 +b^3) a(2a^2 +x ba+ i b^2) + b(y a^2 +j ab +b^2) y = 2 i = 1 x = j x+y = 3, x = 1 i + j = 2, j = 1

OpenStudy (anonymous):

but but I don't see (a+b)(2a^2+ab+b^2) =.= my goal is to just reach that factor

OpenStudy (amistre64):

since the left and right terms need to be the same, we can solve for how we need to split them

OpenStudy (amistre64):

without knowing how to split it, i just used xyij as placeholders to get to: a(2a^2 +x ba+ i b^2) + b(y a^2 +j ab +b^2) now: 2a^2 +x ba+ i b^2 = y a^2 +j ab + b^2 y=2 x=j i=1 , this is the only way itll work for us x+y = 3, and i+j = 2 x+2 = 3, and 1+j = 2 x=1 j=1

OpenStudy (amistre64):

soo, let fill those in: 2a^3 +3ba^2 + 2ab^2 +b^3 2a^3 +(1 ba^2 + 2 ba^2) + (1 ab^2 +1 ab^2) +b^3 (2a^3 +ba^2+ ab^2) + (2ba^2 +ab^2 +b^3) a(2a^2 +ba+ b^2) + b(2a^2 +ab +b^2) (a+b) (2a^2 +ba+ b^2) , as required

OpenStudy (anonymous):

\[2a^3+3ba^2+2ab^2+b^3\] Which methods would you use to factor this? the easiest (I might of picked the hardest lol)

OpenStudy (amistre64):

your pick was fine, otherwise theres a few trial and errors to mess with

OpenStudy (anonymous):

what about one where I don't have to deal with other terms?

OpenStudy (amistre64):

got an example?

OpenStudy (amistre64):

or a method that doesnt deal with new terms ...

OpenStudy (anonymous):

like as how you used x j i y etc

OpenStudy (amistre64):

if we assume (a+b) is a factor, which at first glance it may not be, then division finds the other factor

OpenStudy (amistre64):

2a^2 + ab + b^2 --------------------------- a+b | 2a^3 +3ba^2 + 2ab^2 +b^3 -(2a^3 +2ba^2) ------------ ba^2 +2ab^2 -(a^2b + ab^2) ------------- ab^2 + b^3 -(ab^2 + b^3) ------------ 0

OpenStudy (amistre64):

but there is not way to know if a+b is a factor to start with

OpenStudy (amistre64):

well, there is, let b=-a if it zeros out then (a+b) is a factor

OpenStudy (amistre64):

2a^3 +3(-a)a^2 + 2a(-a)^2 +(-a)^3 2a^3 -3a^3 + 2a^3 - a^3 = 4-4 = 0

OpenStudy (amistre64):

so (a+b) is a factor

OpenStudy (anonymous):

this is way easier than the first part, thanks you.

OpenStudy (anonymous):

wait how did you start with b = -a?

OpenStudy (amistre64):

at this point, there is no 'easier' way to me

OpenStudy (amistre64):

well, if (a+b) is a factor: then f(x) = (a+b)(whatever) is our factorization when (a+b) = 0, then the original f(x) will have to equal 0 a+b = 0 when b=-a, or when a=-b doesnt matter

OpenStudy (amistre64):

knowing that it is a factor, we can do the division to find the (whatever) part

OpenStudy (anonymous):

ooo ok so I can know if those are factors of the function if they are equals to 0

OpenStudy (anonymous):

this would only be easy if we had terms such as a and b right?

OpenStudy (amistre64):

maybe

OpenStudy (anonymous):

ok, I think I got the hang of it. thanks

OpenStudy (amistre64):

good luck :)